OpenStudy (anonymous):

a sphere has a volume of 3600cm3. find the surface area

OpenStudy (anonymous):

use V=(4/3)*pi*r^3 to find the radius, r. then use that r into the surface area forumula: S=4*pi*r^2

OpenStudy (anonymous):

im not sure how to do the \[\sqrt[3]{2700/pi}\]